### v2

# TorchCP: A Python Library for Conformal Prediction

Abstract
Conformal Prediction (CP) has attracted great attention from the research community due to its strict theoretical guarantees. However, researchers and developers still face challenges of applicability and efficiency when applying CP algorithms to deep learning models. In this paper, we introduce TorchCP, a comprehensive PyTorch-based toolkit to strengthen the usability of CP for deep learning models. TorchCP implements a wide range of post-hoc and training methods of conformal prediction for various machine learning tasks, including classification, regression, GNN, and LLM. Moreover, we provide user-friendly interfaces and extensive evaluations to easily integrate CP algorithms into specific tasks. Our TorchCP toolkit, built entirely with PyTorch, enables high-performance GPU acceleration for deep learning models and mini-batch computation on large-scale datasets. With the LGPL license, the code is open-sourced at https://github.com/ml-stat-Sustech/TorchCP and will be continuously updated.
Keywords: Conformal Prediction, Uncertainty Quantification, Machine Learning
1 Introduction
Conformal prediction (CP) is a powerful statistical framework that provides reliable measures of uncertainty for machine learning predictions. With conformal prediction, the model yields prediction sets or intervals designed to contain the ground truths with a specified coverage probability  (Vovk et al., 2005; Shafer and Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos et al., 2020; Manokhin, 2022). In recent years, conformal prediction algorithms have increasingly been integrated into deep learning models, enabling trustworthy predictions for high-stake scenarios in the real world. The trend raises new challenges of applicability and efficiency in implementing conformal prediction algorithms across various tasks of deep learning, such as time series, graph neural networks, and large language models (Zargarbashi et al., 2023; De Toni et al., 2024; Quach et al., 2024; Dey et al., 2024).
Given the increasing popularity of deep learning, it is indispensable to develop an extensive toolkit to enable easy, fast, and flexible integration of CP algorithms in various deep learning applications. There are existing libraries of uncertainty quantification that contain implementations of some CP algorithms, such as Puncc (Mendil et al., 2023), crepes(Boström, 2022), and Fortuna (Detommaso et al., 2024). However, those implementations are typically limited in the scope of application, only supporting classification and regression tasks. It is essential to develop a comprehensive CP library that supports a wider range of deep learning tasks, such as graph node classification, ordinal classification, and text generation with large language models. Besides, most of the current libraries lack support for GPU acceleration, which is crucial for deep learning models, and do not facilitate mini-batch computation for large datasets. This leads to suboptimal computational efficiency and scalability, particularly for large-scale applications. We summarize a feature comparison of current libraries in Table 1.
Table 1: Open-source libraries with support of conformal prediction.
Library Task Support
Classification Regression Graph LLM Training GPU Mini-Batch Backend
nonconformist [^1] https://github.com/donlnz/nonconformist. Numpy
TorchUQ [^2] https://github.com/TorchUQ/torchuq?tab=readme-ov-file. PyTorch
crepes (Boström, 2022) Numpy
Puncc (Mendil et al., 2023) Numpy
MAPIE (Cordier et al., 2023) Numpy
Fortuna (Detommaso et al., 2024) JAX
TorchCP (ours) PyTorch
In this paper, we present TorchCP, a Python library of conformal prediction for deep learning, built on PyTorch with strong GPU acceleration. TorchCP addresses the above limitations through the following crucial features: 1) Comprehensiveness. TorchCP implements a wide range of state-of-the-art CP algorithms, encompassing both post-hoc and training methods. In addition to conventional classification and regression tasks, the framework extends its functionality to various deep learning scenarios, including graph node classification, large language models, and ordinal classification, making it a versatile toolkit for modern machine learning applications. 2) Scalability. TorchCP is designed in a modular fashion and provides user-friendly interfaces for CP algorithms. Researchers can integrate and experiment with various CP algorithms through modular components. The framework’s plug-in architecture also enables the application of CP training algorithms to any existing loss function, offering exceptional flexibility for machine learning practitioners. 3) Efficiency. TorchCP delivers high-performance CP algorithms with strong GPU acceleration. This native PyTorch implementation enables mini-batch computation with high efficiency, providing seamless integration with deep learning models.
In this project, we open-source TorchCP, accompanied by detailed documentation and tutorials, to facilitate the easy, flexible, and efficient implementation of CP algorithms in deep learning applications. We believe that TorchCP will promote the adoption and advancement of conformal prediction within the deep learning community.
Figure 1: The scope of TorchCP.
2 The Architecture Design of TorchCP
2.1 Overview of Implementations
TorchCP implements a wide range of state-of-the-art conformal prediction algorithms, supporting both training approaches and post-hoc methods for various deep learning tasks (Figure 1). Specifically, for classification tasks, TorchCP provides various non-conformity score functions (Gammerman and Vovk, 2007; Löfström et al., 2015; Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2020; Huang et al., 2024), CP algorithms (Vovk, 2012; Tibshirani et al., 2019; Ding et al., 2024), and conformal trainer  (Stutz et al., 2022; Dey et al., 2024; Xi et al., 2024; Liu et al., 2024). For regression tasks, TorchCP implements the standard conformal regression algorithm (Lei et al., 2018), multiple variants of Conformal Quantile Regression (CQR) (Romano et al., 2019; Kivaranovic et al., 2020; Sesia and Candès, 2020), and various conformal algorithms for time series problems (Gibbs and Candes, 2021; Xu and Xie, 2021).
Moreover, TorchCP provides conformal prediction algorithms for node classification of graphs and large language models (LLMs). Specifically, TorchCP implements various variants of APS that are specifically designed for graph structure (Wijegunawardana et al., 2020; Clarkson, 2023; Zargarbashi et al., 2023; Huang et al., 2023; Song et al., 2024), and the certified risk control algorithm for text generation task of LLMs (Quach et al., 2024).
Finally, to facilitate the evaluation and selection of CP methods, TorchCP offers extensive metrics for assessing the marginal/conditional coverage and size efficiency of prediction sets produced by different CP methods. This enables researchers to benchmark the superiority of novel approaches and identify the most suitable existing methods for specific tasks.
2.2 Modular Design
To improve extensibility and flexibility, TorchCP adopts a highly modularized architecture in implementing conformal prediction methods. For each task, TorchCP implements three main components of conformal prediction algorithms: Trainer, Score and Predictor. We illustrate the modular design of TorchCP on the left of Figure 2. The details of the three components are introduced below.
  • Trainer module defines those training algorithms with loss functions for improving conformal prediction, like ConfTr (Stutz et al., 2022) and C-Adapter (Liu et al., 2024).
  • Score module defines those non-conformity scores that quantify how different a new sample is from the training data, e.g., THR (Sadinle et al., 2019) ,APS (Gibbs and Candes, 2021), SAPS (Huang et al., 2024) and CQR (Romano et al., 2019).
  • Predictor module defines the conformal prediction procedure, including the functions of calibration and evaluation. For example, two common-used conformal predictors are Split CP (Vovk et al., 2005) and Class-wise CP (Vovk, 2012). A conformal predictor is defined with a score and a trained model from the trainer. The evaluate function outputs the two measures of predictors: average set size and empirical coverage rate. We provide more metrics for measuring the predictors in utils/metrics.
Example Code for Conformal Prediction. On the right side of Figure 2, we present an example code snippet for multi-class classification. First, we prepare a calibration dataset and a test dataset (Lines 5-6), along with a pre-trained PyTorch model (Lines 9-10). Next, we define a split predictor using the THR score and the trained model (Line 14). The calibration procedure is conducted by running the calibrate function with the cal_dataloader and the significance level $\alpha$ (Line 17). Finally, we obtain the evaluation results through the evaluate function (Lines 20-21). This modular design is compatible with various deep learning tasks, allowing for easy implementation of new conformal prediction algorithms.
Figure 2: An illustration of the TorchCP framework.
3 Conclusion and Future work
This paper presents TorchCP, a Python toolbox for conformal prediction research on deep learning models, built on the PyTorch Library with strong GPU acceleration. TorchCP implements a wide range of state-of-the-art conformal prediction algorithms and provides user-friendly interfaces along with comprehensive evaluation functions. Besides, we provide extensive documents, tutorials, and unit tests (100% test coverage) to improve the usability of TorchCP. The toolbox can support future research and applications of conformal prediction algorithms on deep learning models. We believe that TorchCP library will promote the adoption and advancement of conformal prediction within the deep learning community. In the future, TorchCP will be continuously updated with new algorithms, paradigms, and evaluation protocols. We also welcome any contributions for TorchCP.
Appendix A Details of conformal methods in TorchCP
A.1 Split CP
Conformal prediction (Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth label with a desired probability. Split Conformal Prediction is the most widely-used version of the conformal prediction procedure. Specifically, split conformal prediction divides a given dataset into two disjoint subsets: one for training the base classifier and the other for conformal calibration. We next outline the main process of split conformal prediction:
  1. 1.
    Divide a given dataset into two disjoint subsets: a training fold $\mathcal{D}_{tr}$ and a calibration fold $\mathcal{D}_{cal}$, with $|\mathcal{D}_{cal}|$ being $n$;
  2. 2.
    Train a deep learning model on the training dataset $\mathcal{D}_{tr}$, and define a non-conformity score function $V(\bm{x},y)$;
  3. 3.
    Compute $\widehat{Q}_{1-\alpha}$ as the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ quantile of the calibration scores $\left\{V(\bm{x}_{i},y_{i}):(\bm{x}_{i},y_{i})\in\mathcal{D}_{cal}\right\}$, where $\widehat{Q}_{1-\alpha}$ is defined by
    $\widehat{Q}_{1-\alpha}:=\inf\{Q\in\mathbb{R}:\frac{|\{i:V(\bm{x}_{i},y_{i})\leq Q\}|}{n}\geq\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\};$
  4. 4.
    Use the conformal threshold $\widehat{Q}_{1-\alpha}$ to generate a prediction set for a new instance $\bm{x}_{n+1}$:
    $\mathcal{C}(\bm{x}_{n+1})=\{y:V(\bm{x}_{n+1},y)\leq\widehat{Q}_{1-\alpha}\}.$
A.2 Classification
Here, we consider multi-class classification with $K$ classes. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ to denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ and $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a classification neural network. Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a ($K$-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function. Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$. In the following, we briefly describe methods of conformal prediction supported by TorchCP.
KNN (Gammerman and Vovk, 2007): KNN calculates the score based on the k closest training examples in the feature space. The score is calculated by dividing two sums: the numerator is the sum of the k shortest distances from an input point to other points that share its same label, while the denominator is the sum of the k shortest distances from that input point to points that have different labels than it does.
Margin (Löfström et al., 2015): The margin score is constructed on a well-known concept that a prediction with a larger margin is generally considered to be more conforming. Consequently, the margin score is defined by $V{(\bm{x},y)}:=\pi_{y}-\max_{j\in\mathcal{Y}:{j}\neq y}\pi_{j}.$
THR (Sadinle et al., 2019): The THR measures the similarity between the example and the data space by the conditional probability $P(Y|X)$: $V(\bm{x},y):=1-\pi_{y}(\bm{x})$.
APS (Romano et al., 2020): The score of Adaptive Prediction Sets (APS) calculates the non-conformity score using the cumulative sum of sorted softmax probabilities:
$V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x}),$
where $r_{f}(\bm{x},y)$ denotes the rank of $\pi_{y}(\bm{x})$ among the descending softmax probabilities, and $u$ is an independent random variable satisfying a uniform distribution on $[0,1]$.
RAPS (Angelopoulos et al., 2020): RAPS adds a regularization to penalize noisy tail probabilities and regularize the number of samples in the uncertainty set:
$V(\bm{x},y,u;\pi):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})+\eta\cdot(r_{f}(\bm{x},y)-k_{reg})^{+},$
where $\eta$ represents the weight of regularization, $k_{reg}\geq 0$ are regularization hyper-parameters and $(z)^{+}$ denotes the positive part of $z$.
TOPK (Angelopoulos et al., 2020): TOPK is a score function based solely on the ranking information: $V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u$.
SAPS (Huang et al., 2024): The SAPS method mitigates the issue of probability miscalibration by only retaining the highest probability value and discarding all others:
$V(\bm{x},y,u):=\left\{\begin{array}[]{ll}u\cdot\pi_{max}(\bm{x}),\qquad\text{if\quad}r_{f}(\bm{x},y)=1,\\ \pi_{max}(\bm{x})+(r_{f}(\bm{x},y)-2+u)\cdot\lambda,\quad\text{else},\end{array}\right.$
where $\lambda$ is a hyperparameter representing the weight of ranking information and $\pi_{max}(\bm{x})$ denotes the maximum softmax probability.
Class-wise CP (Vovk, 2012): Class-wise CP partitions the calibration data based on their labels $y$. The non-conformity scores’ quantiles are then computed separately for each class using the corresponding calibration subset.
Clustered CP (Ding et al., 2024): Clustered CP extends Class-wise CP by applying split conformal prediction to class clusters rather than individual classes. It achieves cluster-conditional coverage as an approximation of class-conditional guarantees.
Weighted CP (Tibshirani et al., 2019): Weighted CP addresses distribution shift in Conformal Prediction, where calibration and test data follow different distributions. The core idea is that data points closer to the test point in the feature space provide more valuable information about it.
ConfTr (Stutz et al., 2022): ConfTr simulates the CP process based on end-to-end training by splitting mini-batches into calibration and prediction sets and designing differentiable conformal prediction steps: $\mathcal{L}_{\mathrm{ConfTr}}=\max\left(0,\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})-\kappa\right),$ where $\theta$ is the trainable parameters of deep learning models and $\kappa$ is the target set size. The definition of $\mathcal{C}_{\theta,k}(\bm{x})$ is given by $\sigma((V(\bm{x},k)-\widehat{Q}_{1-\alpha})/\tau)$ where $\tau$ is a temperature hyperparameter and $\sigma$ is the sigmoid function.
ConfTS (Xi et al., 2024): ConfTs is a training method to optimize the $T$ of temperature scaling for improved efficiency of prediction sets. The objective of ConfTS is to minimize the efficiency gap, i.e., $\widehat{Q}_{1-\alpha}-V(\bm{x},y)$. Formally, the loss function of ConfTS is:
$\mathcal{L}_{\mathrm{ConfTS}}=(\widehat{Q}_{1-\alpha}-V(\bm{x},y))^{2}$
C-Adapter (Liu et al., 2024): C-Adapter proposes to optimize an adapter with a Conformal Discriminative (CD) Loss. The CD loss is constructed to improve the $\alpha$-free efficiency of prediction sets: $\mathcal{L}_{\mathrm{CD}}=\frac{1}{K}\sum_{j\in\mathcal{Y}}\{\sigma(V(\bm{x},y)-V(\bm{x},j))\}$.
A.3 Regression
In the context of regression, the goal is to predict a continuous target variable $y\in\mathbb{R}$ based on an input $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, where $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ follows a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Let $f:\mathcal{X}\to\mathcal{F}$ denote a regression model trained on a dataset $\{(\bm{x_{i}},y_{i})\}_{i=1}^{n}$, where $\mathcal{F}$ represents the space of model outputs, $\bm{x}_{i}$ are the features and $y_{i}$ are the observed target values.
ABS (Lei et al., 2018): The ABS score $V(\bm{x},y):=|f(\bm{x})-y|$ is designed to measure the absolute deviation of a prediction $f(\bm{x})$ from the observed target value $y$, where $f(\bm{x})$ is a point estimate of given $\{\bm{x}\}$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha},f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}].$
CQR (Romano et al., 2019): The CQR score $V(\bm{x},y):=\max(f(\bm{x})_{\alpha/2}-y,\;y-$$f(\bm{x})_{1-\alpha/2})$ measures the maximum absolute deviation between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by: $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x})_{\alpha/2}-\widehat{Q}_{1-\alpha},\;f(\bm{x})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}].$
CQRR (Sesia and Candès, 2020): The CQRR score extends the standard CQR method by introducing a scaling mechanism. This score normalizes the deviations between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$ by the width of the interval itself. Formally, for a data pair $(\bm{x},y)$, the CQRR score is defined as $V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha}(\bm{x})}\right),$ where $\hat{s}_{\alpha}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{\alpha/2}+\epsilon$ is the scaling factor, and $\epsilon>0$ is a small constant added to prevent division by zero. The prediction set is given by $\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1})\right].$
CQRM (Sesia and Candès, 2020): The CQRM score extends the CQRR score by introducing two scaling factors to independently normalize the deviations for the lower and upper bounds of the prediction interval. This method requires the model to output predictions at three quantile levels: $[\alpha/2,1/2,1-\alpha/2]$. The two scaling factors are defined as $\hat{s}_{\alpha,\text{lo}}(\bm{x}):=f(\bm{x})_{1/2}-f(\bm{x})_{\alpha/2}+\epsilon$ and $\hat{s}_{\alpha,\text{up}}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{1/2}+\epsilon$. The non-conformity score for a data pair $(\bm{x},y)$ and prediction interval are then defined as:
$V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha,\text{lo}}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha,\text{up}}(\bm{x})}\right).$
Then, the prediction set is given by
$\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{lo}}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{up}}(\bm{x}_{n+1})\right].$
CQRFM (Kivaranovic et al., 2020): The CQRFM score is similar to CQRM, as it also requires the model to output predicted quantiles at $[\alpha/2,1/2,1-\alpha/2]$. However, it normalizes the deviations from the median $f(\bm{x})_{1/2}$ by the fractional distance to the nearest quantile bounds. The two fractional scaling factors are defined in the same way as in CQRM. The non-conformity score and prediction interval are defined as:
$V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{1/2}-y}{\hat{s}_{\alpha,\text{lo}}(\bm{x})},\;\frac{y-f(\bm{x})_{1/2}}{\hat{s}_{\alpha,\text{up}}(\bm{x})}\right).$
Then, the prediction set is given by
$\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{1/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{lo}}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{up}}(\bm{x}_{n+1})\right].$
R2CCP (Guha et al., 2023): R2CCP converts a regression problem into a classification problem by dividing the range space into $K$ bins. The model predicts the probability that each sample belongs to one of these $K$ bins, similar to a standard classification task. To train the regression-to-classification model, R2CCP uses a custom loss function:
$\mathcal{L}(\theta)=\sum_{i=1}^{N}\left(\sum_{k=1}^{K}\left(|y_{i}-m_{k}|^{p}\cdot\pi_{ik}\right)-\tau\cdot\sum_{k=1}^{K}\pi_{ik}\log(\pi_{ik})\right),$
where $\pi_{ik}$ is the predicted probability for the $k$-th bin, $m_{k}$ are the bin midpoints, $\tau$ is a regularization parameter, and $p$ is the exponent applied to the absolute difference between the target value and the bin midpoint. Both $\tau$ and $p$ are hyperparameters. After training the regression-to-classification model, R2CCP defines the non-conformity score as the linear interpolation of the model’s softmax probabilities. Based on these scores, R2CCP outputs the interval where all non-conformity scores are smaller than $\widehat{Q}_{1-\alpha}$.
Ensemble CP (Xu and Xie, 2021): Ensemble CP trains multiple models based on randomly sampled subsets of the data and aggregates their nonconformity scores using a specific aggregation function, such as mean or median. It can update $\widehat{Q}_{1-\alpha}$ over time, making it adaptable to dynamic time series.
ACI (Gibbs and Candes, 2021): ACI dynamically adjusts the confidence level $\alpha_{t}=\alpha_{t-1}+\gamma\left(\alpha-\text{err}_{t}\right)$ based on the historical error rates to adapt to distribution shifts, where $\alpha_{t}$ is the updated confidence level depending on time $t$, $\alpha$ is the target coverage, and $\text{err}_{t}:=\sum_{i=1}^{n}w_{i}\cdot I(y_{i}\notin\hat{C}_{i})$ is the weighted error rate at time step $t$. This dynamic updating enables ACI to maintain reliable prediction sets despite evolving data distributions.
A.4 Graph
Here, we consider node classification tasks. The Graph is represented as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$. Let $\bm{X}:=[\bm{x}_{1},\cdot\cdot\cdot,\bm{x}_{N}]^{T}$ be the node feature matrix, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector for node $v_{i}$. The label of node $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
DAPS (Zargarbashi et al., 2023): DAPS proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily. Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS and RAPS, and $\lambda$ is a diffusion parameter.
SNAPS (Song et al., 2024): SNAPS, which is essentially an enhanced version of the DAPS, aggregates non-conformity scores of nodes with high feature similarity to ego node and one-hop graph structural neighbors. Formally, the score function is shown as: $V(\bm{x}_{i},y,\lambda,\mu):=(1-\lambda-\mu)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{\bm{D}_{s}(i,i)}\sum_{j=1}^{M}\bm{A}_{s}(i,j)\hat{V}(\bm{x}_{j},y)+\frac{\mu}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\bm{A}_{s}$ is the adjacency matrix of $k$-NN graph, which measures the similarity between each node pair, and $\bm{D}_{s}$ is the degree matrix of $\bm{A}_{s}$. $M$ denotes the number of randomly selected nodes as targets used to construct $k$-NN graph. Both $\lambda$ and $\mu$ are hyperparameters, which are used to measure the importance of three parts of non-conformity scores.
NAPS (Clarkson, 2023): NAPS, which adapts the weighted variant of CP from (Barber et al., 2023), assigns a weight to adjacent nodes, and a weight of zero otherwise. NAPS then applies nonexchangeable conformal prediction with the APS scoring function in the following equation: $\mathcal{C}(\bm{x}_{n+1})=\{y\in\mathcal{Y}:V(\bm{x}_{n+1},y)\leq Q_{1-\alpha}\left(\sum_{i}^{n}w_{i}\cdot\delta_{s_{i}}+w_{n+1}\cdot\delta_{+\infty}\right)\},$ where $Q_{\tau}(\cdot)$ denotes the $\tau$-quantile of a distribution and $s_{i}=V(\bm{x}_{i},y_{i})$. Moreover, $w_{i}$ denotes a weight of $s_{i}$ and $\delta_{s_{i}}$ denotes a point mass at $s_{i}$.
CF-GNN (Huang et al., 2023): CF-GNN introduces a topology-aware output correction model, akin to GCN, which employs a conformal-aware inefficiency loss to refine predictions and improve the efficiency of post-hoc CP. For the loss, CF-GNN adopts the optimizing efficiency version of ConfTr.
A.5 LLMs
Conformal LLM (Quach et al., 2024): This paper proposes a novel conformal procedure based on risk control framework (Bates et al., 2021) for language models by reformulating the generation task as a sampling problem rather than enumerating all possible outputs. The core algorithm samples candidate outputs until a calibrated stopping rule is met, while using a rejection rule to filter out low-quality samples. The method guarantees that the output set contains at least one acceptable answer with high probability while maintaining the statistical guarantees of traditional conformal prediction.
Appendix B Benchmark Results
B.1 Classification
In Figure 3, we present the experimental results of split CP on the ImageNet dataset, using THR, APS, RAPS, and SAPS. In Figure 4, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including THR, APS, RAPS, Split CP, Class-wise CP, and Clustered CP. Clustered CP achieves the smallest average size and best class-conditional coverage.
(a) Coverage
(b) Size
Figure 3: Results of image classification with $\alpha=0.1$ on ImageNet dataset.
(a) Coverage
(b) Size
(c) CovGap
Figure 4: Results of image classification with $\alpha=0.1$ on ImageNet dataset .
B.2 Regression
In Figure 5, we present an example of experimental results on the Community dataset (Redmond, 2002), using ABS, CQR, CQRR, CQRM, CQRFM, R2CCP with Split CP.
(a) Coverage
(b) Size
Figure 5: Results of regression with $\alpha=0.1$ on Community dataset.
B.3 Graph Node Classification
In Figure 6, we present an example of experimental results on the CoraML dataset (McCallum et al., 2000), using APS, DAPS, CF-GNN, and SNAPS. The experiments are conducted in a transductive learning setting for node classification tasks.
(a) Coverage
(b) Size
Figure 6: Results of Graph node classification with $\alpha=0.1$ on CoraML dataset.
B.4 LLM
In Table 2, we present an example prediction set, generated by conformal LLM, of the TriviaQA dataset (Joshi et al., 2017). The results show that the generation process of the prediction set stops after sample 5, because the sum of scores collectively exceeds the conformal threshold.
Table 2: Example prediction set of conformal LLM on the TriviaQA dataset.
Question Men Against the Sea and Pitcairn’s Island were two sequels to what famous novel?
Answer Score Label
Sample 1 Mutiny On The Bounty 0.5347 True
Sample 2 Treasure Island 0.5087 False
Sample 3 Treasure Island 0.5087 False
Sample 4 Mutiny on the Bounty 0.1889 True
Sample 5 Robinson Crusoe 0.1384 False
Appendix C Comparison of Computational Efficiency
We evaluated the computational efficiency of TorchCP by comparing its inference time with other existing libraries of conformal prediction. In particular, we focus on the time of calibration and inference on MNIST (LeCun et al., 2010) and ImageNet (Russakovsky et al., 2015) throughout the conformal prediction process. We leave out some libraries that cannot be implemented on the dataset without making significant changes. The results in Figure 7 show that TorchCP is the most efficient implementation among these libraries. TorchCP exhibits superior computational efficiency, attributed to its seamless integration with PyTorch’s architecture and mini-batch parallel processing capabilities.
(a) MNIST
(b) ImageNet
Figure 7: Time comparison of conformal prediction using various open-sourced libraries on MNIST and ImageNet datasets.
Appendix D Extensibility
Each module in TorchCP can be easily extended. In the following, we provide basic guidelines for customizing your own scores, predictors, trainers, and metrics.
D.1 Add new scores
Adding new scores involves two steps:
  1. 1.
    Implementing a New Score Class: new Scores should be implemented in the score module, inheriting from the BaseScore class. Using classification as an example, one could alternatively choose to override the functions: __init__(), __call__(), _calculate_all_label(), and _calculate_single_label().
  2. 2.
    Adding an Interface: After customizing the score class, register it in score/__init__.py.
D.2 Add new predictors
Similar to adding new scores, the addition of new predictors also consists of two steps:
  1. 1.
    Implementing a New Predictor Class: new Predictors should be implemented in the predictor module, inheriting from the BasePredictor class. Using classification as an examples, one could alternatively choose to override the functions: __init__(), calibrate() and predict().
  2. 2.
    Adding an Interface: After customizing the predictor class, register it in predictor/__init__.py.
D.3 Add new trainers
Adding new trainers is similar to steps of D.1 and D.2:
  1. 1.
    Implementing a New Trainer Class: new Trainers should be implemented in the trainer module, inheriting from the BaseTrainer class. Using classification as an example, one could alternatively choose to override the functions: __init__(), train() and validate(). One could pass different loss functions in the __init__() method.
  2. 2.
    Adding an Interface: After customizing the trainer class, register it in trainer/__init__.py.
D.4 Add new metrics
Adding new metrics involves two steps:
  1. 1.
    Implementing a New Metric Function: new Metrics should be implemented in the utils/metrics.py.
  2. 2.
    Adding an Interface: After customizing the metric function, register it by Registry("METRICS") in the utils/metrics.py.