### v3

# TorchCP: A Python Library for Conformal Prediction

Abstract
Conformal prediction (CP) is a robust statistical framework that generates prediction intervals or sets with guaranteed coverage probability, addressing the challenge of quantifying predictive uncertainty in deep learning. Despite advancements in deep learning architectures and datasets, reliable uncertainty estimation remains elusive, making CP increasingly vital. This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning tasks, including classification, regression, graph neural networks, and large language models. TorchCP offers a comprehensive suite of advanced methodologies, a modular design for easy customization, and full GPU-accelerated scalability. Released under the LGPL-3.0 license, TorchCP has gained widespread adoption with over 12,582 PyPi downloads. It is supported by approximately 16,132 lines of code, 564 unit tests achieving 100% coverage, and comprehensive documentation. By bridging statistics and computer science, TorchCP empowers researchers and practitioners to advance conformal prediction in diverse deep learning applications.
Keywords: Conformal Prediction, Deep Learning
1 Introduction
Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets that contain true labels with a specified coverage probability (Vovk et al., 2005). In recent years, deep learning has evolved dramatically with advanced model architectures and vast datasets, yet it struggles to reliably measure predictive uncertainty, making conformal prediction an increasingly attractive solution. Beyond classification and regression, conformal prediction methods have been tailored for diverse deep learning tasks, such as graph neural networks (GNNs) for molecular modeling and large language models (LLMs) for natural language understanding. This underscores the need for a comprehensive and efficient CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across diverse deep learning applications.
This work presents TorchCP, a PyTorch-native library of conformal prediction for deep learning. Firstly, TorchCP implements a wide range of state-of-the-art CP algorithms encompassing both training and inference methods designed for deep learning tasks, e.g., classification (ordinal classification), regression (time series), GNN, and LLM. Secondly, the library’s modular design allows developers to easily use or customize each component (e.g., losses or scores) of the library. Lastly, TorchCP is fully built on PyTorch with GPU support and batch processing, enabling seamless scalability for large-scale deep learning tasks.
The TorchCP library, released under the LGPL-3.0 license, can be installed via PyPi with the command ‘‘pip install torchcp’’ and has been downloaded 12,582 times  [^1] https://pepy.tech/projects/torchcp. The implementation consists of approximately 16,132 lines of code and is thoroughly tested with 564 unit tests, achieving 100% test coverage. We also provide comprehensive online documentation  [^2] https://torchcp.readthedocs.io/en/latest/ to guide users through TorchCP’s features. We hope the library can bridge the gap between statistics and computer science, enabling researchers and practitioners from both backgrounds to advance conformal prediction in deep learning.
2 The library design and implementation
Figure 1: An illustration of the modular design of TorchCP.
2.1 Modular design
For low coupling, we separate the TorchCP library into separate modules that can be easily maintained, tested, and integrated. For each task, the code implementation of the library contains 3 modules: Trainer, Score, and Predictor. Specifically, Trainer defines training algorithms for improving conformal prediction, along with loss functions. Score defines non-conformity score functions that quantify how different a new sample is from the training data, e.g., LAC (Sadinle et al., 2019). Predictor defines CP workflows (e.g., calibration and prediction), such as Split CP  [^3] A detailed introduction of Split CP is presented in Appendix A.1 (Papadopoulos et al., 2002). The modular design allows a straightforward extension to new tasks or new CP methods that arise in this active research field.
We visualize the modular design of TorchCP in Figure 1, highlighting the relationships among the three modules. The Trainer module is used to train or fine-tune deep models, employed by the Predictor module for calibration and prediction. The Score module implements non-conformity score functions, enabling their flexible use across different trainers and predictors. This modular structure empowers users to independently select or customize trainers, score functions, and predictors for diverse conformal prediction tasks, significantly improving flexibility and maintainability. Additionally, TorchCP achieves strong extensibility by providing standardized module interfaces, making it straightforward to define new methods. Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s modular design.
2.2 Implementation
Table 1: The CP algorithms implemented in TorchCP.
Task Component Methods
Classification Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
Score Function LAC (Sadinle et al., 2019), APS (Romano et al., 2020), RAPS (Angelopoulos et al., 2020), SAPS (Huang et al., 2024), KNN (Gammerman and Vovk, 2007), Margin (Löfström et al., 2015), TOPK (Angelopoulos et al., 2020)
Predictor Split CP (Papadopoulos et al., 2002) Class-conditional CP (Vovk, 2012), Clustered CP (Ding et al., 2024), Weighted CP (Tibshirani et al., 2019), RC3P (Shi et al., 2024)
Regression Score Function ABS (Papadopoulos et al., 2002), NORABS (Vovk et al., 2005), CQR (Romano et al., 2019), CQRR (Sesia and Candès, 2020), CQRM (Sesia and Candès, 2020), CQRFM (Kivaranovic et al., 2020), R2CCP (Guha et al., 2023)
Predictor Ensemble CP (Xu and Xie, 2021), ACI (Gibbs and Candes, 2021), AgACI (Zaffran et al., 2022)
GNN Trainer&Loss CF-GNN (Huang et al., 2023)
Score Function DAPS (Zargarbashi et al., 2023), SNAPS (Song et al., 2024)
Predictor NAPS (Clarkson, 2023)
LLM Predictor Conformal LLM (Quach et al., 2024)
Methodological scope
Table 1 outlines the CP algorithms included in TorchCP across 4 deep learning tasks: classification, regression (time series), GNN, and LLM , along with their corresponding evaluation metrics in Table 4. In total, TorchCP provides 6 training algorithms, 16 score functions, and 10 predictors tailored to various tasks. A brief overview of all supported methods and their advantages is available in Appendix A.
Batch Processing and GPU Support
TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities. It enables efficient computation on large datasets and scales seamlessly to real-world deep learning applications. Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
Tutorials
TorchCP aims to provide a user-friendly framework for conformal prediction in deep learning. The library contains illustrative code showcasing functionality for common tasks (See the “examples” folder  [^4] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples), and Appendix C offers a practical guide to help users identify relevant examples. In Appendix D, we provide reproduced results of classical CP methods, confirming the correctness of our implementation. Below, we present a generic example  [^5] The full code of the example is available at https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_conftr_cifar100.py to show the usability of TorchCP. First, we employ ConfTr (Stutz et al., 2022) to train a model for conformal prediction (optional, using standard trained models also work):
1from torchcp.classification.trainer import ConfTrTrainer
2trainer = ConfTrTrainer(init_model, alpha=0.1) # Create a trainer with a NN
3trained_model = trainer.train(train_loader, num_epochs=10) # Training 10 epochs
With the trained model, users can proceed with conformal prediction by selecting an appropriate score function and predictor. As an example, we implement Split CP with a miscoverage rate of $\alpha=0.1$ using the LAC score :
1predictor = SplitPredictor(score_function=LAC(), model=trained_model, alpha=0.1)
2predictor.calibrate(cal_loader) # Compute the threshold on calibrate data
3predict_set = predictor.predict(test_instance) # Predict the set for an instance
3 Related library
Conformal Prediction (CP) is a distribution-free framework for quantifying the uncertainty of machine learning model predictions. While some uncertainty quantification libraries, such as PUNCC (Mendil et al., 2023), implement certain CP algorithms, their limited efficiency, scalability, and algorithmic coverage restrict their use in deep learning. For instance, most libraries support only a narrow range of tasks, such as regression and classification, and lack support for complex applications like generative modeling or time-series forecasting. Additionally, the absence of GPU acceleration and batch processing in these libraries makes them inefficient for large-scale scenarios. In contrast, TorchCP offers a clean interface, enhanced extensibility, high efficiency, and a comprehensive set of CP algorithms. For example, TorchCP reduces inference time by up to 90% compared to PUNCC on large-scale datasets (see Appendix D.2 for details). We summarize the comparison of different libraries in Table 2 and discuss the advantages of TorchCP ’s modular design in Appendix E.
Table 2: Feature comparison of open-source libraries supporting conformal prediction. “Training” refers to CP-specific training algorithms, and “Online” denotes online CP.
Library Task Support Functionality Support Backend
Classification Regression GNN LLM Training Online GPU Batch
nonconformist (Linusson et al., 2020) Numpy
TorchUQ (Zhao et al., 2024) PyTorch
Crepes (Boström, 2022) Numpy
PUNCC (Mendil et al., 2023) Numpy
MAPIE (Cordier et al., 2023) Numpy
Fortuna (Detommaso et al., 2024) JAX
TorchCP (ours) PyTorch
4 Conclusion and Future work
In this work, we introduced TorchCP, a PyTorch-native library that seamlessly integrates conformal prediction (CP) into deep learning workflows. Supporting a wide range of state-of-the-art CP algorithms, TorchCP leverages PyTorch’s GPU acceleration and batch processing, offering a modular design for easy customization and user-friendly documentation for accessibility. Looking ahead, we plan to expand TorchCP by incorporating emerging CP algorithms for evolving architectures like diffusion models and multimodal models, while optimizing for real-time performance. Through open-source collaboration on GitHub, we aim to foster community-driven innovation, empowering reliable AI in fields such as healthcare, finance, and autonomous systems.
Appendix A Details of Conformal Prediction Methods in TorchCP
In this section, we provide brief introductions to the methods implemented in TorchCP. Furthermore, Table 3 offers practical guidelines to help users select appropriate methods based on their specific application needs.
Table 3: Advantages or use cases of CP methods across different tasks.
Component Method Advantage/Use case
Classification
Score LAC To achieve the smallest prediction sets
APS To approximate conditional coverage
RAPS For smaller, more stable sets than APS
TOPK A simple baseline with only logit ranking
SAPS Improving efficiency, conveying instance-wise uncertainty
Margin A non-conformity function based on prediction margin
KNN A score function based solely on feature-space nearest neighbors without a predictive model.
Predictor Class-conditional CP To achieve coverage for every class
Clustered CP For class-conditional coverage with limited data per class
Weighted CP Conformal prediction under covariate shift
Loss & Trainer ConfTr Efficient prediction sets with conformal-aware loss
C-Adapter Optimizing efficiency based on $\alpha$-free loss
ConfTS Improving efficiency via temperature scaling
Regression
Score ABS Simple and interpretable residual-based scores
NORABS Improving robustness of prediction intervals via noise-aware normalization
CQR Flexible interval width via quantile regression
R2CCP Shorter prediction intervals than CQR via recasting regression as classification
Predictor ACI Reliable coverage under distributional shifts
AgACI Aggregates ACI predictions with adaptive weighting
Ensemble CP Stable prediction intervals without data splitting via model ensembling
GNN
Score DAPS Improves efficiency of APS using structural neighbors
SNAPS Extends DAPS with feature neighbors for better efficiency
NAPS Constructing prediction sets for inductive node classification via Weighted CP
Loss & Trainer CF-GNN Reducing set size by attaching an adapter with a topology-aware inefficiency loss
LLM
Predictor Conformal LLM To achieve statistically guaranteed prediction sets for language models through calibrated sampling
A.1 Split CP
Conformal prediction (Papadopoulos et al., 2002; Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth labels with a desired probability. Split Conformal Prediction (Papadopoulos et al., 2002) (Split CP) is the most widely-used version of the conformal prediction procedure. Specifically, Split CP divides a given dataset into two disjoint subsets: one for training the base model and the other for conformal calibration. We next outline the main process of split conformal prediction:
  1. 1.
    Divide a given dataset into two disjoint subsets: a training fold $\mathcal{D}_{tr}$ and a calibration fold $\mathcal{D}_{cal}$, with $|\mathcal{D}_{cal}|$ being $n$;
  2. 2.
    Train a model on the training dataset $\mathcal{D}_{tr}$, and then define a non-conformity score function $V(\bm{x},y)$;
  3. 3.
    Compute $\widehat{Q}_{1-\alpha}$ as the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ quantile of the calibration scores $\left\{V(\bm{x}_{i},y_{i}):(\bm{x}_{i},y_{i})\in\mathcal{D}_{cal}\right\}$, where $\widehat{Q}_{1-\alpha}$ is defined by
    $\widehat{Q}_{1-\alpha}:=\inf\{Q\in\mathbb{R}:\frac{|\{i:V(\bm{x}_{i},y_{i})\leq Q\}|}{n}\geq\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\};$
  4. 4.
    Use the conformal threshold $\widehat{Q}_{1-\alpha}$ to generate a prediction set for a new instance $\bm{x}_{n+1}$:
    $\mathcal{C}(\bm{x}_{n+1})=\{y:V(\bm{x}_{n+1},y)\leq\widehat{Q}_{1-\alpha}\}.$
A.2 Classification
Here, we provide a brief overview of conformal prediction methods for multi-class classification tasks as supported by TorchCP. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ to denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ and $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a classification neural network. Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a ($K$-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function. Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$.
LAC (Sadinle et al., 2019): The LAC measures the similarity between the example and the data space by the conditional probability $P(Y|X)$: $V(\bm{x},y):=1-\pi_{y}(\bm{x})$.
APS (Romano et al., 2020): The score of Adaptive Prediction Sets (APS) calculates the non-conformity score using the cumulative sum of sorted softmax probabilities:
$V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x}),$
where $r_{f}(\bm{x},y)$ denotes the rank of $\pi_{y}(\bm{x})$ among the descending softmax probabilities, and $u$ is an independent random variable satisfying a uniform distribution on $[0,1]$.
RAPS (Angelopoulos et al., 2020): RAPS adds a regularization to penalize noisy tail probabilities and regularize the number of samples in the uncertainty set:
$V(\bm{x},y,u;\pi):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})+\eta\cdot(r_{f}(\bm{x},y)-k_{reg})^{+},$
where $\eta$ represents the weight of regularization, $k_{reg}\geq 0$ are regularization hyper-parameters and $(z)^{+}$ denotes the positive part of $z$.
TOPK (Angelopoulos et al., 2020): TOPK is a score function based solely on the ranking information: $V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u$.
SAPS (Huang et al., 2024): The SAPS method mitigates the issue of probability miscalibration by only retaining the highest probability value and discarding all others:
$V(\bm{x},y,u):=\left\{\begin{array}[]{ll}u\cdot\pi_{max}(\bm{x}),\qquad\text{if\quad}r_{f}(\bm{x},y)=1,\\ \pi_{max}(\bm{x})+(r_{f}(\bm{x},y)-2+u)\cdot\lambda,\quad\text{else},\end{array}\right.$
where $\lambda$ is a hyperparameter representing the weight of ranking information and $\pi_{max}(\bm{x})$ denotes the maximum softmax probability.
KNN (Gammerman and Vovk, 2007): KNN calculates the score based on the k closest training examples in the feature space. The score is calculated by dividing two sums: the numerator is the sum of the k shortest distances from an input point to other points that share its same label, while the denominator is the sum of the k shortest distances from that input point to points that have different labels than it does.
Margin (Löfström et al., 2015): The margin score is constructed on a well-known concept that a prediction with a larger margin is generally considered to be more conforming. Consequently, the margin score is defined by $V{(\bm{x},y)}:=\pi_{y}-\max_{j\in\mathcal{Y}:{j}\neq y}\pi_{j}.$
Class-conditional CP (Vovk, 2012): Class-conditional CP partitions the calibration data based on their labels $y$. The non-conformity scores’ quantiles are then computed separately for each class using the corresponding calibration subset.
Clustered CP (Ding et al., 2024): Clustered CP extends Class-conditional CP by applying split conformal prediction to class clusters rather than individual classes. It achieves cluster-conditional coverage as an approximation of class-conditional guarantees.
Weighted CP (Tibshirani et al., 2019): Weighted CP addresses distribution shift in Conformal Prediction, where calibration and test data follow different distributions. The core idea is that data points closer to the test point in the feature space provide more valuable information about it.
RC3P (Shi et al., 2024): RC3P addresses the inefficiency of class-conditional conformal prediction, which often results in large prediction sets, especially for imbalanced classification tasks. The core idea is to selectively perform class-wise thresholding only for labels with low top-k error, by integrating a label rank calibration step into standard class-wise conformal calibration.
ConfTr (Stutz et al., 2022): ConfTr simulates the CP process based on end-to-end training by splitting mini-batches into calibration and prediction sets and designing differentiable conformal prediction steps: $\mathcal{L}_{\mathrm{ConfTr}}=\max\left(0,\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})-\kappa\right),$ where $\theta$ is the trainable parameters of deep learning models and $\kappa$ is the target set size. The definition of $\mathcal{C}_{\theta,y}(\bm{x})$ is given by $\sigma((\widehat{Q}_{1-\alpha}-V(\bm{x},y))/\tau)$ where $\tau$ is a temperature hyperparameter and $\sigma$ is the sigmoid function.
ConfTS (Xi et al., 2025): ConfTs is a training method to optimize the $T$ of temperature scaling for improved efficiency of prediction sets. The objective of ConfTS is to minimize the efficiency gap, i.e., $\widehat{Q}_{1-\alpha}-V(\bm{x},y)$. Formally, the loss function of ConfTS is:
$\mathcal{L}_{\mathrm{ConfTS}}=(\widehat{Q}_{1-\alpha}-V(\bm{x},y))^{2}$
C-Adapter (Liu et al., 2024): C-Adapter proposes to optimize an adapter with a Conformal Discriminative (CD) Loss. The CD loss is constructed to improve the $\alpha$-free efficiency of prediction sets: $\mathcal{L}_{\mathrm{CD}}=\frac{1}{K}\sum_{j\in\mathcal{Y}}\{\sigma(V(\bm{x},y)-V(\bm{x},j))\}$.
SCPO (Bellotti, 2021): SCPO proposes an approach to train the deep learning model directly by optimizing for maximum predictive efficiency. To enable end-to-end training, the conformal predictor is approximated by a differentiable surrogate objective. Gradient descent is then used to minimize the following loss:
$\mathcal{L}_{\mathrm{SCPO}}=\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})+\lambda\left(\mathcal{C}_{\theta,y}(\bm{x})-(1-\alpha)\right)^{2},$
where $\lambda>0$ controls the relative importance of the two objective components.
Uncertainty-aware (UA) Loss (Einbinder et al., 2022): Uncertainty-aware Loss introduces a novel training objective that integrates conformal calibration directly into model training by encouraging the conformity scores to follow a uniform distribution. The novel uncertainty-aware component of this loss is:
$\mathcal{L}_{u}=\sup_{w\in[0,1]}\left|\mathbb{E}[\mathds{1}(V(\bm{x},y,u)-w]\right|,$
where $V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})$.
A.3 Regression
In the context of regression, the goal is to predict a continuous target variable $y\in\mathbb{R}$ based on an input $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, where $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ follows a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Let $f:\mathcal{X}\to\mathcal{F}$ denote a regression model trained on a dataset $\{(\bm{x_{i}},y_{i})\}_{i=1}^{n}$, where $\mathcal{F}$ represents the space of model outputs, $\bm{x}_{i}$ are the features and $y_{i}$ are the observed target values.
ABS (Papadopoulos et al., 2002): The ABS score $V(\bm{x},y):=|f(\bm{x})-y|$ is designed to measure the absolute deviation of a prediction $f(\bm{x})$ from the observed target value $y$, where $f(\bm{x})$ is a point estimate of given $\{\bm{x}\}$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha},f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}].$
NORABS (Vovk et al., 2005): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\hat{\sigma}(\bm{x})}$ is a normalized variant of ABS. It uses models $f(\bm{x})$ and $\hat{\sigma}(\bm{x})$ trained to maximize the likelihood of the data with respect to $\mathbb{E}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]$ and $\sqrt{\text{Var}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]}$, respectively. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1}),f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1})].$
CQR (Romano et al., 2019): The CQR score $V(\bm{x},y):=\max(f(\bm{x})_{\alpha/2}-y,\;y-$$f(\bm{x})_{1-\alpha/2})$ measures the maximum absolute deviation between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by: $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x})_{\alpha/2}-\widehat{Q}_{1-\alpha},\;f(\bm{x})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}].$
CQRR (Sesia and Candès, 2020): The CQRR score extends the standard CQR method by introducing a scaling mechanism. This score normalizes the deviations between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$ by the width of the interval itself. Formally, for a data pair $(\bm{x},y)$, the CQRR score is defined as $V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha}(\bm{x})}\right),$ where $\hat{s}_{\alpha}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{\alpha/2}+\epsilon$ is the scaling factor, and $\epsilon>0$ is a small constant added to prevent division by zero. The prediction set is given by $\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1})\right].$
CQRM (Sesia and Candès, 2020): The CQRM score extends the CQRR score by introducing two scaling factors to independently normalize the deviations for the lower and upper bounds of the prediction interval. This method requires the model to output predictions at three quantile levels: $[\alpha/2,1/2,1-\alpha/2]$. The two scaling factors are defined as $\hat{s}_{\alpha,\text{lo}}(\bm{x}):=f(\bm{x})_{1/2}-f(\bm{x})_{\alpha/2}+\epsilon$ and $\hat{s}_{\alpha,\text{up}}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{1/2}+\epsilon$. The non-conformity score for a data pair $(\bm{x},y)$ and prediction interval are then defined as:
$V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha,\text{lo}}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha,\text{up}}(\bm{x})}\right).$
Then, the prediction set is given by
$\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{lo}}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{up}}(\bm{x}_{n+1})\right].$
CQRFM (Kivaranovic et al., 2020): The CQRFM score is similar to CQRM, as it also requires the model to output predicted quantiles at $[\alpha/2,1/2,1-\alpha/2]$. However, it normalizes the deviations from the median $f(\bm{x})_{1/2}$ by the fractional distance to the nearest quantile bounds. The two fractional scaling factors are defined in the same way as in CQRM. The non-conformity score and prediction interval are defined as:
$V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{1/2}-y}{\hat{s}_{\alpha,\text{lo}}(\bm{x})},\;\frac{y-f(\bm{x})_{1/2}}{\hat{s}_{\alpha,\text{up}}(\bm{x})}\right).$
Then, the prediction set is given by
$\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{1/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{lo}}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{up}}(\bm{x}_{n+1})\right].$
R2CCP (Guha et al., 2023): R2CCP converts a regression problem into a classification problem by dividing the range space into $K$ bins. The model predicts the probability that each sample belongs to one of these $K$ bins, similar to a standard classification task. To train the regression-to-classification model, R2CCP uses a custom loss function:
$\mathcal{L}(\theta)=\sum_{i=1}^{N}\left(\sum_{k=1}^{K}\left(|y_{i}-m_{k}|^{p}\cdot\pi_{ik}\right)-\tau\cdot\sum_{k=1}^{K}\pi_{ik}\log(\pi_{ik})\right),$
where $\pi_{ik}$ is the predicted probability for the $k$-th bin, $m_{k}$ are the bin midpoints, $\tau$ is a regularization parameter, and $p$ is the exponent applied to the absolute difference between the target value and the bin midpoint. Both $\tau$ and $p$ are hyperparameters. After training the regression-to-classification model, R2CCP defines the non-conformity score as the linear interpolation of the model’s softmax probabilities. Based on these scores, R2CCP outputs the interval where all non-conformity scores are smaller than $\widehat{Q}_{1-\alpha}$.
Ensemble CP (Xu and Xie, 2021): Ensemble CP trains multiple models based on randomly sampled subsets of the data and aggregates their nonconformity scores using a specific aggregation function, such as mean or median. It can update $\widehat{Q}_{1-\alpha}$ over time, making it adaptable to dynamic time series.
ACI (Gibbs and Candes, 2021): ACI dynamically adjusts the confidence level $\alpha_{t}=\alpha_{t-1}+\gamma\left(\alpha-\text{err}_{t}\right)$ based on the historical error rates to adapt to distribution shifts, where $\alpha_{t}$ is the updated confidence level depending on time $t$, $\alpha$ is the target coverage, and $\text{err}_{t}:=\sum_{i=1}^{n}w_{i}\cdot I(y_{i}\notin\hat{C}_{i})$ is the weighted error rate at time step $t$. This dynamic updating enables ACI to maintain reliable prediction sets despite evolving data distributions.
AgACI (Zaffran et al., 2022): AgACI is a parameter-free conformal prediction method for time series that builds on ACI by using online expert aggregation. It runs multiple models in parallel with different learning rates $\gamma_{k}$ and adaptively aggregates their prediction intervals using weighted averages, where weights are based on the historical pinball loss. This approach eliminates the need to tune $\gamma$ manually and ensures valid coverage while improving interval efficiency, especially under temporal dependencies.
A.4 GNN
We consider the prediction tasks on graph data. The Graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$. Let $\bm{X}:=[\bm{x}_{1},\cdot\cdot\cdot,\bm{x}_{N}]^{T}$ be the node feature matrix, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector for node $v_{i}$. The label of $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
DAPS (Zargarbashi et al., 2023): DAPS proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily. Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS, and $\lambda$ is a diffusion parameter.
SNAPS (Song et al., 2024): SNAPS, which is essentially an enhanced version of the DAPS, aggregates non-conformity scores of nodes with high feature similarity to ego node and one-hop graph structural neighbors. Formally, the score function is shown as: $V(\bm{x}_{i},y,\lambda,\mu):=(1-\lambda-\mu)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{\bm{D}_{s}(i,i)}\sum_{j=1}^{M}\bm{A}_{s}(i,j)\hat{V}(\bm{x}_{j},y)+\frac{\mu}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\bm{A}_{s}$ is the adjacency matrix of $k$-NN graph, which measures the similarity between each node pair, and $\bm{D}_{s}$ is the degree matrix of $\bm{A}_{s}$. $M$ denotes the number of randomly selected nodes as targets used to construct $k$-NN graph. Both $\lambda$ and $\mu$ are hyperparameters, which are used to measure the importance of three parts of non-conformity scores.
NAPS (Clarkson, 2023): NAPS, which adapts the weighted variant of CP from (Barber et al., 2023), assigns a weight to adjacent nodes, and a weight of zero otherwise. NAPS then applies nonexchangeable conformal prediction with the APS scoring function in the following equation: $\mathcal{C}(\bm{x}_{n+1})=\{y\in\mathcal{Y}:V(\bm{x}_{n+1},y)\leq Q_{1-\alpha}\left(\sum_{i}^{n}w_{i}\cdot\delta_{s_{i}}+w_{n+1}\cdot\delta_{+\infty}\right)\},$ where $Q_{\tau}(\cdot)$ denotes the $\tau$-quantile of a distribution and $s_{i}=V(\bm{x}_{i},y_{i})$. Moreover, $w_{i}$ denotes a weight of $s_{i}$ and $\delta_{s_{i}}$ denotes a point mass at $s_{i}$.
CF-GNN (Huang et al., 2023): CF-GNN introduces a topology-aware output correction model, akin to Graph Convolutional Network, which employs a conformal-aware inefficiency loss to refine predictions and improve the efficiency of post-hoc CP. For the loss, CF-GNN adopts the optimizing efficiency version of ConfTr.
A.5 LLM
Conformal LLM (Quach et al., 2024): This paper proposes a novel conformal procedure based on a risk control framework (Bates et al., 2021) for language models by reformulating the generation task as a sampling problem rather than enumerating all possible outputs. The core algorithm samples candidate outputs until a calibrated stopping rule is met, while using a rejection rule to filter out low-quality samples. The method guarantees that the output set contains at least one acceptable answer with high probability while maintaining the statistical guarantees of traditional conformal prediction.
A.6 Metrics
In this section, we provide a brief overview of the evaluation metrics supported by TorchCP for various predictive tasks, as summarized in Table 4. These metrics facilitate rigorous and fine-grained assessment of prediction sets, offering deep insights into model uncertainty.
Table 4: Implemented evaluation metrics in TorchCP.
Task Metrics
Classification Coverage, Size, CovGap (Ding et al., 2024), WSC (Romano et al., 2020), SSCV (Angelopoulos et al., 2020), Violated classes (Kasa and Taylor, 2023), DiffViolation (Angelopoulos et al., 2020), , etc.
Regression Coverage, Size
GNN Coverage, Size, Singleton ratio (Zargarbashi et al., 2023)
LLM SetLoss, SSCL (Quach et al., 2024)
Appendix B Extensibility
Each module in TorchCP can be easily extended. In the following, we provide basic guidelines for customizing your own scores, predictors, trainers, and metrics.
B.1 Add new scores
Adding new scores involves two steps:
  1. 1.
    Implementing a New Score Class: new Scores should be implemented in the score module, inheriting from the BaseScore class. Using classification as an example, one could alternatively choose to override the functions: __init__(), __call__(), _calculate_all_label(), and _calculate_single_label().
  2. 2.
    Adding an Interface: After customizing the score class, register it in score/__init__.py.
B.2 Add new predictors
Similar to adding new scores, the addition of new predictors also consists of two steps:
  1. 1.
    Implementing a New Predictor Class: new Predictors should be implemented in the predictor module, inheriting from the BasePredictor class. Using classification as an example, one could alternatively choose to override the functions: __init__(), calibrate(), and predict().
  2. 2.
    Adding an Interface: After customizing the predictor class, register it in predictor/__init__.py.
B.3 Add new trainers
Adding new trainers is similar to steps of B.1 and B.2:
  1. 1.
    Implementing a New Trainer Class: new Trainers should be implemented in the trainer module, inheriting from the BaseTrainer class. Using classification as an example, one could alternatively choose to override the functions: __init__(), train() and validate(). One could pass different loss functions in the __init__() method.
  2. 2.
    Adding an Interface: After customizing the trainer class, register it in trainer/__init__.py.
B.4 Add new metrics
Adding new metrics involves two steps:
  1. 1.
    Implementing a New Metric Function: New metrics should be implemented in the utils/metrics.py.
  2. 2.
    Adding an Interface: After customizing the metric function, register it by Registry(‘‘METRICS’’) in the utils/metrics.py.
Appendix C Example Codes for TorchCP
TorchCP provides a collection of example codes in the “examples" folder  [^6] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples of the official codebase, designed to help users quickly get started with various conformal prediction tasks. In this section, we briefly introduce representative examples for the four supported tasks.
C.1 Example codes for offline conformal prediction
This part provides brief descriptions of example codes under the offline setting:
  1. 1.
    Classification. This example code [^7] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100.py applies Split CP with the LAC score on CIFAR-100 using a pretrained ResNet-20. It evaluates the coverage and average set size of the resulting prediction sets.
  2. 2.
    Regression. The example code of regression [^8] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/regression_cqr_synthetic.py shows the complete workflow of applying Split CP with the Conformalized Quantile Regression score (Romano et al., 2019) on a synthetic dataset in regression task, including model training, calibration, prediction interval generation, and evaluation.
  3. 3.
    GNN. The node classification example [^9] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/gnn_transductive_coraml.py implements a Graph Convolutional Network (GCN) (Kipf and Welling, 2017) for node classification on the CoraML dataset (McCallum et al., 2000) under a transductive setting. Conformal prediction is then applied to generate prediction sets with guaranteed coverage.
  4. 4.
    LLM. The LLM example [^10] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/llm_ConformalLM_TriviaQA.py demonstrates how to deploy conformal prediction in large language model generation tasks, i.e., TriviaQA (Joshi et al., 2017).
C.2 Example codes for online conformal prediction
Online conformal prediction means that after predicting a test point, its true label is immediately revealed, and the test sample is incorporated into the calibration set. TorchCP supports this setting as well, with the following examples:
  1. 1.
    Regression. In TorchCP, the online setup of the regression task is integrated into ACIPredictor (Gibbs and Candes, 2021). Therefore, we added an example code  [^11] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/timeseries_aci_synthetic.py for the regression task in TorchCP.
  2. 2.
    Classification. For classification tasks, this online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data. To clarify this, we have added an online conformal prediction example 121212 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
Appendix D Empirical Results
This section presents empirical results of TorchCP. We first reproduce several classical conformal prediction methods across various tasks. Then, we benchmark the computational efficiency of TorchCP against several representative conformal prediction libraries.
D.1 Evaluation results
Classification.
In Figure 2, we present the experimental results of Split CP on ImageNet, using LAC, APS, RAPS (penalty=0.1, kreg =0), and SAPS (weight=0.2). In Figure 3, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including Split CP, Class-conditional CP, and Clustered CP with APS. The number of trials is 1000.
(a) Coverage
(b) Size
Figure 2: Results of image classification with $\alpha=0.1$ on ImageNet dataset.
(a) Coverage
(b) Size
(c) CovGap
Figure 3: Conditional coverage results on ImageNet image classification with $\alpha=0.1$.
Regression.
In Figure 4, we present an example of experimental results on the Community dataset (Redmond, 2002), using ABS, CQR, CQRR, CQRM, CQRFM, R2CCP with Split CP. The number of trials is 1000.
(a) Coverage
(b) Size
Figure 4: Results of regression with $\alpha=0.1$ on Community dataset.
GNN.
In Figure 5, we present experimental results on the CoraML dataset (McCallum et al., 2000) for node classification under a transductive learning setting. We evaluate four methods: APS, DAPS, CF-GNN, and SNAPS, each conducted over 1000 independent trials. For DAPS, the diffusion parameter is set to 0.5. In CF-GNN, the topology-aware correction model is implemented using a two-layer GCN. For SNAPS, the hyperparameters are set as follows: $\lambda=\frac{1}{3}$, $\mu=\frac{1}{3}$, and $k=20$. These hyperparameter choices follow the default or recommended configurations reported in the respective original papers.
(a) Coverage
(b) Size
Figure 5: Results of Graph node classification with $\alpha=0.1$ on CoraML dataset.
LLM.
In Table 5, we present an example prediction set, generated by conformal LLM, of the TriviaQA dataset (Joshi et al., 2017). The generation process terminates after the fifth sample, as the cumulative score surpasses the conformal threshold.
Table 5: An Example of prediction sets from Conformal LLM on TriviaQA.
Question Men Against the Sea and Pitcairn’s Island were two sequels to what famous novel?
Answer Score Label
Answer 1 Mutiny On The Bounty 0.5347 True
Answer 2 Treasure Island 0.5087 False
Answer 3 Treasure Island 0.5087 False
Answer 4 Mutiny on the Bounty 0.1889 True
Answer 5 Robinson Crusoe 0.1384 False
D.2 Comparison of computational efficiency
D.2.1 Classification Task
We evaluated the computational efficiency of TorchCP in image classification tasks by comparing its calibration and inference time with several existing CP libraries. Specifically, we use two standard benchmarks: MNIST (LeCun et al., 2010) and ImageNet (Russakovsky et al., 2015). Throughout all experiments, we used Split CP with APS as the conformal prediction method. Moreover, TorchCP supports two data processing modes: full-batch (processing the entire dataset at once), and batch (processing data in mini-batches). In our experiments, we set the batch size to 128 for batch processing. We use NVIDIA H100 for GPU computation and AMD EPYC 9654 for CPU computation.
MNIST. We compared the computational efficiency of five libraries: PUNCC, MAPIE, Fortuna, Crepes, and TorchCP (GPU-based full-batch processing). In the dataset, we use 48,000 samples for training, 6,000 for calibration, and 6,000 for testing. For TorchCP, we use a LeNet-5 model for training and inference using PyTorch on a GPU. For PUNCC, MAPIE, and Crepes, we obtain model predictions using PyTorch on a GPU, and then convert output probabilities to NumPy arrays for conformal calibration and inference using CPU. For Fortuna, we train the LeNet-5 model using the built-in training algorithm and perform model predictions with JAX on a GPU. We only report the total runtime for Fortuna as it ties the conformal calibration and inference processes together. Figure 6(a) presents the running times of the five libraries on the MINST dataset. The results show that TorchCP achieves the lowest calibration and inference time among the five libraries. Specifically, TorchCP completes calibration and inference for the entire dataset in just 0.07 seconds, using only 3.6% of the runtime needed by PUNCC (1.93 seconds). Moreover, the Crepes library exhibits the longest inference time on the MNIST dataset, likely due to its computation of p-values for each test sample’s non-conformity scores by comparing them against the entire calibration set. This substantial improvement highlights the efficiency advantage of TorchCP on small-scale datasets.
ImageNet. We compared the computational efficiency of four libraries: PUNCC, MAPIE, Crepes, and TorchCP (GPU-based batch processing). In the dataset, we use a pre-trained ResNet-50 model from torchvision and select 25,000 images from the validation set for calibration and another 25,000 for testing. For TorchCP, we leverage GPU-based batch processing to efficiently run the conformal pipeline. For PUNCC, MAPIE, and Crepes, we conducted experiments on the CPU, as they lack batch processing support, and their GPU-based full-batch processing causes out-of-memory errors. Figure 6(b) shows the results of running times of three libraries on the ImageNet dataset. The results demonstrate that TorchCP achieves the shortest calibration and inference times, cutting approximately 90% of the time required by PUNCC (104s vs. 995s). This significant enhancement underscores the scalability of TorchCP on large-scale datasets and models.
(a) MNIST
(b) ImageNet
Figure 6: Runtime comparison of different conformal prediction libraries.
Choices of Batch processing. To assess the efficiency of two processing modes (Full-batch and batch), we conduct experiments with TorchCP on both MNIST and ImageNet, using the same data splits and models as in prior experiments. We utilize GPU-accelerated computing for all experiments except the full-batch mode on ImageNet, where the dataset’s size exceeds GPU memory capacity. Figures 7(a) and  7(b) present the running times of different data processing modes on MNIST and ImageNet, respectively. On the small-scale dataset - MNIST, using full-batch processing on GPU achieves the fastest implementation among the three choices, requiring only 26% computational time of that with CPU. When it comes to the ImageNet dataset, GPU-based batch processing completes calibration and inference in approximately 104.29 seconds, compared to 1022.55 seconds for CPU-based processing, achieving a 9.8× speedup. These findings suggest that GPU-based full-batch processing is optimal for small-scale datasets where GPU memory is sufficient to load the whole dataset. Conversely, GPU-based batch processing is preferable for large-scale datasets like ImageNet, offering significant runtime advantages and enabling scalable, efficient deployment.
(a) MNIST
(b) ImageNet
Figure 7: Runtime comparison of different data processing modes.
D.2.2 Regression Task
We evaluate the computational efficiency and predictive performance of TorchCP and an R-based package, AdaptiveConformal (Susmann et al., 2023) on a time series regression task, using the France electricity forecasting dataset [^13] https://cran.r-project.org/web/packages/opera/index.html. Specifically, we use the first 350 time steps to train a SARIMA model with order (1, 0, 0) and seasonal order (1, 1, 0), and construct prediction intervals for the remaining time steps. To assess computational efficiency, we apply the ACI algorithm(Gibbs and Candes, 2021) using both TorchCP and AdaptiveConformal. Both AdaptiveConformal and TorchCP are executed on CPU in this experiment to ensure a fair comparison. Moreover, since ACI performs recalibration for each test instance, the reported time includes both calibration and inference.
Figure 8(a) reports the runtime results of these two packages. The results show that TorchCP achieves significantly higher computational efficiency compared to the R-based package, reducing running time from 6.1 seconds to around 0.06 seconds. Figures 8(b) and 8(c) illustrate the prediction intervals produced by AdaptiveConformal and TorchCP, respectively. The two sets of intervals are nearly identical, demonstrating the correctness and consistency of the TorchCP implementation. This substantial improvement highlights the efficiency advantage of TorchCP for time series regression.
(a) Time
(b) AdaptiveConformal
(c) TorchCP
Figure 8: Comparison of time consumption and prediction intervals between AdaptiveConformal and TorchCP on the France electricity dataset.
Appendix E The Advantage of our modular design
The modular design of TorchCP prioritizes low coupling to ensure flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks. In contrast, previous packages (e.g., PUNCC and Fortuna) focus on a monolithic design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP). This limits extensibility when new CP predictors or custom score functions are needed. TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components. In what follows, we elaborate on the concrete benefits of our modular design:
  • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows. For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement scores for predictors.
  • Modification isolation: our module design ensures that changes to one module (e.g., refining the APS score) do not affect other modules. This isolation minimizes the risk of unintended side effects, reducing maintenance complexity and preventing errors from propagating across the system. For instance, tightly coupled systems like PUNCC require redefining the predictor to modify a scoring function, while TorchCP confines changes to the target module, streamlining updates and preserving stability.
  • Research extensibility: by adhering to standardized interfaces, new functionality (e.g., a novel non-conformity score) can be integrated into TorchCP without altering existing modules. This contrasts with PUNCC, where introducing new scoring methods demands extensive predictor modifications. Extensibility supports rapid iteration and long-term evolvability, enabling efficient adaptation to new research or applications.
  • Ease of testing and debugging: decoupled modules can be tested and debugged independently, as their interactions are governed by well-defined interfaces. This modularity simplifies identifying and resolving issues within a single module (e.g., debugging a Trainer algorithm) without navigating complex dependencies. In TorchCP, this reduces testing overhead and accelerates development cycles compared to tightly coupled systems, where debugging often requires analyzing the entire pipeline.